OpenStudy (seb.cal):

Please help me. I got this question wrong on my test and it won't tell me why. Solve for x: 1 over 2 plus 4 over 2 x equals quantity x plus 4 over 10

OpenStudy (seb.cal):

@bbypanda16 hey, do you know how to do this? or do you have the correct answer? Algebra 2, 05.08, solving Rational equations.

OpenStudy (seb.cal):

@ChantySquirrel1129**

OpenStudy (bbypanda16):

Yes let me figure it out first :)

OpenStudy (chantysquirrel1129**):

Ummm, it's a little hard to see how it's written Dx

OpenStudy (chantysquirrel1129**):

\[\frac{ 1 }{ 2 }+\frac{ 4 }{ 2 }x=x+\frac{ 4 }{ 10 }\] ?

OpenStudy (bbypanda16):

1 over 2 plus 4 over 2 x equals quantity x plus 4 over 10 x=-1 over 10

OpenStudy (bbypanda16):

@seb.cal

OpenStudy (seb.cal):

@bbypanda16 Sorry, I wasn't specific on the answer choices.. A. x = -4 B. x = 5 C. Both A and B. D. No Solution. \[\frac{ 1 }{ 2 } + \frac{ 4 }{ 2x } = \frac{ x+4 }{ 10 }\] @ChantySquirrel1129**

OpenStudy (seb.cal):

@Azbadar Please help me with this...

OpenStudy (seb.cal):

@fomf

OpenStudy (seb.cal):

@~dumbass~

OpenStudy (seb.cal):

@Itxhel

OpenStudy (anonymous):

@itxhel

OpenStudy (anonymous):

@fomf

OpenStudy (anonymous):

@matthew120

OpenStudy (bbypanda16):

x=5 :)

OpenStudy (bbypanda16):

@seb.cal

OpenStudy (seb.cal):

tysm! ♥ @bbypanda16

OpenStudy (bbypanda16):

Noooo problem xD <3